Solve the following spherical triangle using Law of Sine and Cosine

A= 137°28´, C=135°44´, c= 160° Find : a


1
Expert's answer
2022-05-26T23:53:57-0400

"A = 137\degree28\rq = 137.47\degree"

"С = 135\degree44\rq = 135.73\degree"

Using Law of Sine and Cosine for spherical triangle "\frac{sin (A)}{sin (a)} = \frac{sin (B)}{sin (b)} = \frac{sin (C)}{sin (c)}",

hence "\frac{sin (A)}{sin (a)} = \frac{sin (C)}{sin (c)}".


According to the task "\frac{sin (137.47\degree)}{sin (a)} = \frac{sin (135.73\degree)}{sin (160\degree)}" .


Then "sin (a) = \frac{sin (137.47\degree) * sin (160\degree)}{sin (135.73\degree)};"

"sin (a) = \frac{0.68 * 0.34 }{0.7};"

"sin (a) = 0.33"


Therefore "a = 19.3\degree".

However "19.3\degree < 90\degree", so the sine ambiguity appears (if angle "A > 90\degree" (given), then side a should be "> 90\degree").


That is why "a = 180\degree - 19.3\degree = 160.7\degree = 160\degree42\rq."


Answer: "160\degree42\rq"